OpenStudy (anonymous):

How do you evaluate: secx cscx= 2cscx

OpenStudy (bahrom7893):

Secx = 1/Cosx Cscx = 1/Sinx

OpenStudy (bahrom7893):

1/Cosx * 1/Sinx = 1/(SinxCosx) = (SinxCosx)/(SinxCosx)^2

OpenStudy (anonymous):

i am trying to solve the equation.

OpenStudy (bahrom7893):

oh you want to find x?

OpenStudy (anonymous):

yes! sorry!

OpenStudy (bahrom7893):

@.Sam. i'm stuck

OpenStudy (bahrom7893):

Sec(x)Csc(x)=2Csc(x) Sec(x)Csc(x)-2Csc(x) = 0 Csc(x) (Sec(x)-2) = 0 Csc(x) = 0 Sec(x) = 2 Idk that's weird. I don't see a soln here.

sam (.sam.):

lol

sam (.sam.):

wait..

sam (.sam.):

I think the answer should be \[2\csc(2x)\]

sam (.sam.):

I meant the RHS

OpenStudy (anonymous):

can you do this: secx(csc^2x-2) and then set each equal to 0 and get x= pi/2, 3pi/2 and x= pi/4,3pi/4,5pi/4 and 7pi/4